I need help solving the problem 2^(x^2)=7^x, or 2 to the x-squared power equals 7 to the x power. Thanks in advance for your help.
Answer by lwsshak3(11628)

I need help solving the problem 2^(x^2)=7^x, or 2 to the x-squared power equals 7 to the x power.
**
2^(x^2)=7^x
take log of both sides
x^2log2=xlog7
x^2log2-xlog7=0
x(xlog2-log7)=0
x=0
or
xlog2-log7=0
xlog2=log7
x=log7/log2
x≈2.8074